OpenStudy (anonymous):

write an equation for the specific line: through (4,-12) and parallel to 4x+3y=12

OpenStudy (anonymous):

4x+3y=-20

OpenStudy (aroub):

parallel=same slope ,y=mx+b m= -4/3x , (4,-12) -12=-4/3(4) +b -12=-16/3+b -36/3+16/3=b -23/3=b so y=-4/3x-23/3

OpenStudy (anonymous):

solve 4x + 3y = 12 for y... so that it is in (y = mx + b) form... 4x + 3y = 12 minus 4x from each side and get: 3y = -4x +12 divide by 3 on both sides and that will be y = (-4/3)x + 4 which is the y = mx + b form... now you can see what m is... so now you know the slope is (-4/3) = m So you can use the point slope formula which is y - y1 = m(x -x1) you know the point is (4 -12) and you know m is (-4/3) so the equation is using point slope is y - (-12) = (-4/3)(x - 4) now solve... first distribute the (-4/3) y+12 = (-4/3)x + (16/3) {two neg. = pos} subtract 12 to both sides y = (-4/3)x + (16/3) - 12 y = (-4/3)x - 20/3 unless I made an error somewhere. foo! that was fun.